OpenStudy (anonymous):

this is a little long, sorry. A printer charges $2.80 for binding + 10 cents per page for each copy of a community group's annual report. Write the cost equation & answer the following questions. 1. what is the fixed cost? $ 2. what is the variable cost? $ 3 let y represent the cost of the annual report and x represent the # of pages. what is the cost equation? y= 4. what will it cost to produce a 28 page report 5. if the report cost was $9.80, the report is ____ pages.

OpenStudy (anonymous):

1. $2.80 2. $0.10 3. y=0.10x+2.80 4. y=0.10(28)+2.80=5.60 5. 9.80=0.10x+2.80 x=70